# Nudging Nutrition: Lessons from the Danish “Fat Tax”

### v2

@@ C.1 Bacon
Figure 23: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (1) of Tables 8 to 10.
C.2 Butter
@@
Figure 24: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (2) of Tables 8 to 10.
C.3 Cheese
@@
Figure 25: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (3) of Tables 8 to 10.
C.4 Cream
@@
Figure 26: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (4) of Tables 8 to 10.
C.5 Liver sausage
@@
Figure 27: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (5) of Tables 8 to 10.
C.6 Margarine
@@
Figure 28: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (6) of Tables 8 to 10.
C.7 Salami
@@
Figure 29: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (7) of Tables 8 to 10.
C.8 Sour cream
@@
Figure 30: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (8) of Tables 8 to 10.
@@
Appendix D Robustness checks
D.1 Results when excluding Southern-Danish households
@@
(1) (2) (3) (4) (5) (6) (7) (8) Bacon Butter Cheese Cream Liver sausage Margarine Salami Sour cream Panel A: Quarterly total weight consumed in gr/ml Danish fat tax (ATT) -51.848 56.607 $-417.038^{**}$ 6.714 -116.145 118.680 -58.025 18.472 (se) (39.344) (87.142) (122.812) (74.980) (51.767) (98.194) (37.582) (40.566) % change -9.39% 3.72% -15.42% 0.50% -10.99% 6.93% -11.39% 3.30% obs. 23,844 34,792 26,029 33,376 25,232 32,761 25,186 30,678 nr. of Danish hh. 1,690 1,892 1,841 1,700 1,813 1,647 1,752 1,520 nr. of North.-German hh. 553 1,616 698 1,596 627 1,614 660 1,409 Panel B: Quarterly total expenditure in Euro cents Danish fat tax (ATT) -25.626 $227.535^{***}$ $-294.203^{*}$ $63.645^{*}$75.550 $79.356^{**}$42.121 22.480 (se) (28.542) (41.428) (110.695) (22.569) (31.786) (23.360) (37.458) (11.112) % change -7.13% 24.07% -12.52% 13.16% -13.10% 15.28% -7.78% 13.42% obs. 23,844 34,792 26,029 33,376 25,232 32,761 25,186 30,678 nr. of Danish hh. 1,690 1,892 1,841 1,700 1,813 1,647 1,752 1,520 nr. of North.-German hh. 553 1,616 698 1,596 627 1,614 660 1,409 Panel C: Quarterly total amount (in packages) consumed Danish fat tax (ATT) -0.058 0.278 -0.370 -0.272 -0.438 0.406 0.005 0.014 (se) (0.164) (0.334) (0.383) (0.262) (0.205) (0.215) (0.207) (0.148) % change -3.09% 4.64% -8.43% -8.41% -13.95% 11.19% 0.19% 1.13% obs. 23,844 34,792 26,029 33,376 25,232 32,761 25,186 30,678 nr. of Danish hh. 1690 1,892 1,841 1,700 1,813 1,647 1,752 1,520 nr. of North.-German hh. 553 1,616 698 1,596 627 1,614 660 1,409 Significance level: * $p<0.1$* $p<0.05$** $p<0.01$. Standard errors are clustered at the household level. % change compares the estimated tax average to the pre-tax sample mean for Danish households. The estimates are obtained using the doubly robust DID approach, controlling for age of head of household, number of children under 15, gender, ISCED, income level, and household size.
D.2 Results under unconditional parallel trends for the weight outcome
Figure 31: Bacon.
Figure 32: Butter.
Figure 33: Cheese.
Figure 34: Cream.
Figure 35: Liver sausage.
Figure 36: Margarine.
Figure 37: Salami.
@@ D.2 Results under unconditional parallel trends for the weight outcome
Figure 39: Average quarterly weight consumed in gr/ml for the different product categories. The figures depict the trends when unconditional parallel trends is assumed and provide insights into the strength of the influence of the household characteristics on the differences in trends between Northern-German and Danish households. The trends are similar to those observed under conditional parallel trends (see Appendix C), indicating that the household characteristics are not strongly influencing differences in consumption trends between Danish and Northern-German households.
Appendix E Supplemental material for the heterogeneity analysis
(a) Average quarterly weight consumed (gr/ml).
(b) Average quarterly expenditure in Euro cents.
(c) Average quarterly number of packages.
Figure 40: Doubly robust DID estimates of the causal effect of the Danish fat tax during the post-tax period (January 2013 to December 2014) compared to the last taxed quarter (October 2012 to December 2012) for low, medium and high-income households. Low-income households are defined as households with very low income or low income. High-income households are defined as households with very high income or high income. 95% confidence intervals in red.

### v3

@@ D.1 Bacon
Figure 23: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (1) of Tables 12 to 14.
D.2 Butter
@@
Figure 24: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (2) of Tables 12 to 14.
D.3 Cheese
@@
Figure 25: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (3) of Tables 12 to 14.
D.4 Cream
@@
Figure 26: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (4) of Tables 12 to 14.
D.5 Liver sausage
@@
Figure 27: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (5) of Tables 12 to 14.
D.6 Margarine
@@
Figure 28: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (6) of Tables 12 to 14.
D.7 Salami
@@
Figure 29: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (7) of Tables 12 to 14.
D.8 Sour cream
@@
Figure 30: Time average treatment effects, $ATT(q)$. The two vertical dashed lines indicate the introduction and abolishment of the Danish fat tax, respectively. Red estimates and simultaneous 95% confidence intervals for pre-tax periods, and blue estimates and simultaneous 95% confidence intervals for tax and post-tax periods, clustered at the household level. Under conditional parallel trends, red estimates should be close to zero. Estimates are based on the doubly robust Equation (2). Point estimates and standard errors are reported in Column (8) of Tables 12 to 14.
@@
Appendix E Robustness checks
E.1 Results when excluding Southern-Danish households
@@
(1) (2) (3) (4) (5) (6) (7) (8) Bacon Butter Cheese Cream Liver sausage Margarine Salami Sour cream Panel A: Quarterly total weight consumed in gr/ml Danish fat tax (ATT) -62.619 56.607 $-411.433^{**}$ 6.714 $-108.259^{*}$ 118.680 -49.521 18.472 (se) (34.487) (88.085) (130.721) (77.347) (39.793) (98.689) (35.699) (39.050) p-value 0.220 0.671 0.030 0.952 0.055 0.413 0.337 0.746 % change -11.34% 3.72% -15.78% 0.50% -10.23% 6.93% -13.64% 3.30% obs. 23,073 34,792 25,939 33,376 23,399 32,761 24,857 30,678 nr. of Danish hh. 1,690 1,892 1,824 1,700 1,813 1,647 1,706 1,520 nr. of North.-German hh. 480 1,616 698 1,596 428 1,614 658 1,409 Panel B: Quarterly total expenditure in Euro cents Danish fat tax (ATT) -29.556 $227.535^{***}$ $-298.029^{*}$ $63.645^{*}$ $-48.729^{*}$ $79.356^{**}$45.785 22.480 (se) (22.454) (41.715) (104.902) (23.633) (20.870) (23.583) (38.443) (10.042) p-value 0.355 0.000 0.051 0.071 0.098 0.021 0.407 0.120 % change -8.22% 24.07% -13.07% 13.16% -8.44% 15.28% -10.44% 13.42% obs. 23,073 34,792 25,939 33,376 23,399 32,761 24,857 30,678 nr. of Danish hh. 1,690 1,892 1,824 1,700 1,813 1,647 1,706 1,520 nr. of North.-German hh. 480 1,616 698 1,596 428 1,614 658 1,409 Panel C: Quarterly total amount (in packages) consumed Danish fat tax (ATT) -0.117 0.278 -0.349 -0.272 $-0.321^{*}$ 0.406 -0.004 0.014 (se) (0.142) (0.331) (0.409) (0.274) (0.121) (0.208) (0.213) (0.146) p-value 0.567 0.555 0.561 0.494 0.071 0.179 0.991 0.949 % change -6.17% 4.64% -8.45% -8.41% -10.23% 11.19% -0.20% 1.13% obs. 23,073 34,792 25,939 33,376 23,399 32,761 24,857 30,678 nr. of Danish hh. 1,690 1,892 1,824 1,700 1,813 1,647 1,706 1,520 nr. of North.-German hh. 480 1,616 698 1,596 428 1,614 658 1,409 Significance level: * $p<0.1$* $p<0.05$** $p<0.01$. Standard errors are clustered at the household level. % change compares the estimated tax average to the pre-tax sample mean for Danish households. The estimates are obtained using the doubly robust DID approach, controlling for age of head of household, number of children under 15, gender, ISCED, income level, and household size.
E.2 Results under unconditional parallel trends for the weight outcome
@@
Figure 39: Average quarterly weight consumed in gr/ml for the different product categories. The figures depict the trends when unconditional parallel trends is assumed and provide insights into the strength of the influence of the household characteristics on the differences in trends between Northern-German and Danish households. The trends are similar to those observed under conditional parallel trends (see Appendix D), indicating that the household characteristics are not strongly influencing differences in consumption trends between Danish and Northern-German households.
E.3 Results under unconditional parallel trends for the total expenditure outcome
Figure 40: Bacon.
Figure 41: Butter.
Figure 42: Cheese.
Figure 43: Cream.
Figure 44: Liver sausage.
Figure 45: Margarine.
Figure 46: Salami.
Figure 47: Sour cream.
Figure 48: Average expenditure (Euro cents) for the different product categories. The figures depict the trends when unconditional parallel trends is assumed and provide insights into the strength of the influence of the household characteristics on the differences in trends between Northern-German and Danish households. The trends are similar to those observed under conditional parallel trends (see Appendix D), indicating that the household characteristics are not strongly influencing differences in consumption trends between Danish and Northern-German households.